OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9 7 −7 11 −11

OpenStudy (anonymous):

x=-11

OpenStudy (anonymous):

5x +2 = 4x - 9 Your equal sign is a wall.. to jump the wall, change your sign so +2 becomes -2 on both sides 5x +2 - 2 = 4x -9-2 5x=4x-11 Jump the wall again with the 4x 5x-4x = 4x-4x-11 1x = -11